if f(x) = { x/tan2x, x>0 , x+2, x≥ 0

find the type of discontinuity and state whether it is continuous



1
Expert's answer
2021-12-16T17:03:46-0500

Assume given function as:

f(x) = { x/tan2x, x<0 , x+2, x≥ 0

Solution:

"LHL=\lim_{x\rightarrow0^-}\dfrac x{\tan 2x}
\\=\lim_{h\rightarrow0}\dfrac {0-h}{\tan 2(0-h)}
\\=\lim_{h\rightarrow0}\dfrac {-h}{-\tan 2h}
\\=\lim_{h\rightarrow0}\dfrac {h}{\tan 2h}
\\=\lim_{h\rightarrow0}\dfrac {1}{\dfrac{\tan 2h}{2h}\times 2}
\\=\dfrac {1}{1\times 2} [\because \lim_{x\rightarrow0}\dfrac{\tan x}{x}=1]
\\=\dfrac 12"

"RHL=\lim_{x\rightarrow0^+} (x+2)
\\=\lim_{h\rightarrow0} (0+h+2)
\\=\lim_{h\rightarrow0}\ 2
\\=2"

"\because LHL\ne RHL"

"\\\therefore" Given function is discontinuous at "x=0" which is jump discontinuity.